# Automorphisms of free metabelian Lie algebras

### v2

@@ 5. Tame automorphisms
Every matrix of the form (11) belongs to $T$ if $\mathrm{ldeg}(a)\geq 2$.
Proof. The proof practically repeats the proof of Theorem 1 from [39].
Due to the additivity of (11) with respect to $a$, we can assume that $a=y_{1}^{i_{1}}y_{2}^{i_{2}}\ldots y_{n}^{i_{n}}$. If $i_{1}=0$ then $I+aF_{231}\in T$ by the definition of $T$. Assume that $i_{1}\geq 1$.
Case 1. Suppose that $i_{j}\geq 1$ for some $2\leq j\leq n-1$. Set
$A=I+y_{1}^{i_{1}}\ldots y_{j}^{i_{j}-1}F_{23n},\ \ B=I-y_{j+1}^{i_{j+1}}\ldots y_{n}^{i_{n}}F_{jn1}.$
A direct calculation using the multiplication tables gives that
$[A,B]=ABA^{-1}B^{-1}=I+aF_{231}\in T$
since $A,B\in T$.
Case 2. Let $i_{2}=i_{3}=\ldots=i_{n-1}=0$ and $i_{n}\neq 0$. Then $a=y_{1}^{i_{1}}y_{n}^{i_{n}}$ and we get
$aF_{231}=y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1}-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1}$
by the Jacobi identity. If $i_{n}\geq 2$ then
$I+y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1},\ I-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1}\in T$
by Case 1. Consequently,
$I+aF_{231}=(I+y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}F_{2n1})(I-y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}F_{3n1})\in T.$
If $i_{n}=1$, then set $\alpha=I+E_{42}$. Notice that $I+y_{1}^{i_{1}}y_{2}F_{231}\in T$ by Case 1. Hence
$(I+E_{42})(I+y_{1}^{i_{1}}y_{2}F_{231})^{\alpha}(I-E_{42})=I+y_{1}^{i_{1}}y_{2}F_{231}+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})-y_{1}^{i_{1}}y_{4}F_{341}$
$=(I+y_{1}^{i_{1}}y_{2}F_{231})(I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))(I-y_{1}^{i_{1}}y_{4}F_{341})\in T.$
Since the first and the last factors of this product belong to $T$ it follows that
$C=I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})\in T.$
The Jacobi identity implies that
$y_{4}F_{231}=y_{3}F_{241}-y_{2}F_{341}.$
Consequently,
$3aF_{231}=y_{1}^{i_{1}}[(y_{4}F_{231}-y_{2}F_{341})+(y_{4}F_{231}+y_{3}F_{241})]$
and, by additivity,
$I+3aF_{231}=[I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341})][I+y_{1}^{i_{1}}(y_{4}F_{231}+y_{3}F_{241})].$
The first factor is $C$ and belongs to $T$. The second factor also belongs to $T$ since
$(23)(C^{-1})^{(23)}(23)=(23)((I+y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))^{-1})^{(23)}(23)$
$=(23)(I-y_{1}^{i_{1}}(y_{4}F_{231}-y_{2}F_{341}))^{(23)}(23)=I+(y_{4}F_{231}+y_{3}F_{241}),$
where $(23)=(x_{1},x_{3},x_{2},x_{4},\ldots,x_{n})$. Consequently, $I+3aF_{231}\in T$. If the characteristic of $K$ is not $3$, then its conjugation by $(x_{1},3x_{2},x_{3},\ldots,x_{n})$ gives that $I+aF_{231}\in T$.
If $n\geq 5$, then conjugating (11) by $(4n)$, we can make $i_{4}>0$. Case 1 gives that $I+aF_{231}\in T$.
Case 3. Let $i_{2}=i_{3}=\ldots=i_{n-1}=i_{n}=0$. In this case $a=y_{1}^{i_{1}}$ and $s=i_{1}\geq 2$ since $\mathrm{ldeg}(a)\geq 2$. Case 2 gives that $I+y_{4}y_{1}^{s-1}F_{231}\in T$. Set $\beta=I+E_{14}\in GL_{n}(K)$. Using the multiplication tables we get
$\beta(I+y_{4}y_{1}^{s-1}F_{231})^{\beta}\beta^{-1}=I+(y_{4}+y_{1})y_{1}^{s-1}F_{231}-(y_{4}+y_{1})y_{1}^{s-1}F_{234}$
$=(I+(y_{4}+y_{1})y_{1}^{s-1}F_{231})(I-(y_{4}+y_{1})y_{1}^{s-1}F_{234})$
$=(I+y_{4}y_{1}^{s-1}F_{231})(I+y_{1}^{s}F_{231})(I-y_{4}y_{1}^{s-1}F_{234})(I-y_{1}^{s}F_{234})\in T.$
Since all factors of this product except $I+y_{1}^{s}F_{231}$ belong to $T$, this implies that $I+y_{1}^{s}F_{231}$ also belongs to $T$. $\Box$
Here is a reformulation of this lemma in the language of automorphisms.
Corollary 3.
Every Chein automorphism of the form (5) with $\mathrm{ldeg}(f)\geq 4$ is tame.
Proof. By Corollary 1, any such Chein automorphism is a product of linear conjugates of Chein automorphisms of the form (7) with $\mathrm{ldeg}(a)\geq 2$. Then $J(\tau)$ is a matrix of the form (11) and belongs to $T$ by Lemma 5. $\Box$
This corollary contradicts the results of [42].
Corollary 4.
Assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$. Then the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) and a cubic automorphism (2).
Proof. The group of all tame automorphisms $\mathrm{TAut}(M_{n})$ is generated by all linear automorphisms, as well as a quadratic automorphism (1) [39]. All Chein automorphisms of the form (7) with $\mathrm{ldeg}(a)\geq 2$ belong to $\mathrm{TAut}(M_{n})$. By the additivity of automorphisms of the form (7) with respect to $a$, we may assume that $a$ is homogeneous of degree $\leq 1$. If $a$ does not depend on $y_{1}$ then (7). Again, by additivity, we may assume that $a=y_{1}$. In this case (7) is the cubic automorphism (2). $\Box$
Lemma 6.
Every matrix of the form
$A(a)=I+ay_{1}F_{121}+ay_{2}F_{122}$
belongs to $T$ if $\mathrm{ldeg}(a)\geq 2$.
Proof. By Lemma 5,
$B=I-by_{1}F_{341}-by_{2}F_{342}=(I-by_{1}F_{341})(I-by_{2}F_{342})\in T$
if $\mathrm{ldeg}(b)\geq 1$. Set $C=I+F_{123}\in T$. Then
$[B,C]=BCB^{-1}C^{-1}=I+by_{1}y_{4}F_{121}+by_{2}y_{4}F_{122}=A(y_{4}b)\in T.$
Consequently, the statement of the lemma is true if $y_{4}$ divides $a$. Obviously, the same is true if $y_{i}$ divides $a$ for all $3\leq i\leq n$. By the additivity of $A(a)$ with respect to $a$, we may assume that $a\in K[y_{1},y_{2}]$.
Set $\alpha=I+E_{14}\in T$. If $b\in K[y_{1},y_{2}]$ then, using the multiplication tables, we get
$\alpha A(y_{4}b)^{\alpha}\alpha^{-1}=I+y_{1}(y_{4}+y_{1})bF_{121}+y_{2}(y_{4}+y_{1})bF_{122}-y_{2}(y_{4}+y_{1})bF_{124}$
$=A(y_{4}b)A(y_{1}b)(I-y_{2}(y_{4}+y_{1})bF_{124})\in T.$
Since $A(y_{4}b)\in T$ and $I-y_{2}(y_{4}+y_{1})bF_{124}\in T$ by Lemma 5, this implies that $A(y_{1}b)\in T$. Therefore the statement of the lemma is true if $a\in K[y_{1},y_{2}]$ and $y_{1}$ divides $a$. Obviously, the same is true for $y_{2}$. $\Box$
Corollary 5.
@@
Proof. By Lemma 6 the Jacobian matrices of these automorphisms belong to $T$. $\Box$
Corollary 6.
Every exponential automorphism of lower degree $\geq 5$ is tame.
Proof. Let
$\exp(\mathrm{ad}(m))=(x_{1}+[m,x_{1}],x_{2}+[m,x_{2}],\ldots,x_{n}+[m,x_{n}])$
be an exponential automorphism with $\mathrm{ldeg}(m)\geq 4$. By the additivity with respect to $m$ we may assume that $m=[x_{1},x_{2}]a$ with $\mathrm{ldeg}(a)\geq 2$. Then the Jacobian matrix of $\exp(\mathrm{ad}(m))$ is
$I+\sum_{i}y_{i}aF_{12i}=(I+y_{1}aF_{121}+y_{2}aF_{122})\prod_{i\geq 3}(I+y_{i}aF_{12i}).$
This matrix belongs to $T$ by Lemmas 5 and 6. $\Box$
This corollary contradicts the main result of [6].
6. Almost tame automorphisms
For the remainder of this article, we always assume that $n\geq 4$ and $K$ is an arbitrary field. By Corollary 2, every matrix of the form (11) and its conjugates belong to $S$. For this reason we do not need the proof of Lemma 5, where the condition that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$ was used. The other proofs of Section 5 did not use this condition.
Lemma 7.
Every matrix of the form
$C(a)=I+ay_{1}F_{121}+ay_{2}F_{122}$
with $\mathrm{ldeg}(a)\geq 1$ belongs to $S$.
@@ 6. Almost tame automorphisms
$B=I-by_{1}F_{341}-by_{2}F_{342}=(I-by_{1}F_{341})(I-by_{2}F_{342})\in S.$
Repeating the proof of Lemma 6 with this data, we get that $C(a)\in S$ for all $a$ with $\mathrm{ldeg}(a)\geq 1$. $\Box$
Notice that Lemma 7 does not include only the matrix
(12) $I+y_{1}F_{121}+y_{2}F_{122}$
of the same type, which is the Jacobian matrix of (3).
@@
Proof. Using the decomposition of the Jacobian matrix of an exponential automorphism from the proof of Corollary 6, we conclude that it belongs to $S$ by Lemma 7 if its lower degree $\geq 3$. Consequently, every exponential automorphism of lower degree $\geq 4$ belongs to $\mathrm{ATAut}(M_{n})$. $\Box$
7. Special matrices
We use matrices instead of matrix notation in many places because it makes the calculations easier to understand. To some extent this is due to the application of Suslin’s theorem [61].
The main aim of this section is to prove that every matrix of the form
(13) $I+\begin{bmatrix}y_{n}C&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}*&*&\ldots&*\end{matrix}&0\end{bmatrix}\in G$
belongs to $S$. We will provide several lemmas to the proof of this result. A matrix of the form (13) will be called special.
@@ 7. Special matrices
Let $B$ be any matrix from $SL_{n-1}(U,y_{n}U)$. Then $B=I_{n-1}+y_{n}C$ for some matrix $C\in\mathrm{Mat}_{n-1}(U)$ and there exists a unique special matrix $A$ such that
$A=I+\left[\begin{array}[]{cc}y_{n}C&0\\ &0\\ \end{array}\right].$
Proof. Obviously, $B=I_{n-1}+y_{n}C$ for some matrix $C\in\mathrm{Mat}_{n-1}(U)$. If $C=[c_{ij}]$ then
$A=I+\begin{bmatrix}y_{n}D&\begin{matrix}0\\ .\\ .\\ .\\ 0\end{matrix}\\ \begin{matrix}-\sum_{i=1}^{n-1}y_{i}d_{i1}&-\sum_{i=1}^{n-1}y_{i}d_{i2}&\ldots&-\sum_{i=1}^{n-1}y_{i}d_{i(n-1)}\end{matrix}&0\end{bmatrix}$
belongs to $G$ by Lemma 2. If $C=0$ then $A=I$ by Lemma 2. This implies the uniqueness of $A$. $\Box$
If $B\in SL_{n}(U)$ is a matrix of the form $B=\begin{bmatrix}C&0\\ 0&I_{n-k}\end{bmatrix}$, where $C\in SL_{k}(U)$ for some $1\leq k\leq n$, then we say that $B$ is $k$-reducible and $B_{k}=C$ will be called its $k$-reduction. Calculations with $k$-reducible matrices from $G$ can be done by using their $k$-reductions.
Set $A_{n-1}=B$ for matrices $A,B$ satisfying the conditions of Lemma 8. We use this notation for special matrices despite the fact that $A$ is not n1n-1-reducible in general. But Lemma 16 allows us to use the matrix $B$ instead of $C$ in some calculations. Moreover, if $B$ is $k$-reducible for some $1\leq k\leq n-1$, then $B_{k}$ will also be called the $k$-reduction of the special matrix $A$ and the special matrix $A$ will be called special $k$-reducible. Let us specify how can we use $k$-reductions of special matrices. Using (6) it is easy to check the following statements:
$(1)$ If $A$ and $B$ are two special $k$-reducible matrices, then $AB$ is a special $k$-reducible and $(AB)_{k}=A_{k}B_{k}$;
$(2)$ If $A$ is special $k$-reducible and $B\in SL_{n}(U)$ (not necessarily $B\in G$) is $k$-reducible, then $A^{B}=BAB^{-1}$ is special $k$-reducible and $(A^{B})_{k}=B_{k}A_{k}B_{k}^{-1}$;
$(3)$ If $A$ is special $k$-reducible and $\alpha\in SL_{n}(K)$ is $k$-reducible, then the linear conjugation $\alpha A^{\alpha}\alpha^{-1}$ of $A$ by $\alpha$ is special $k$-reducible and $(\alpha A^{\alpha}\alpha^{-1})_{k}=\alpha_{k}A_{k}^{\alpha_{k}}\alpha_{k}^{-1}$.
Lemma 9.
Every special matrix of the form
$A(h,g)=I+hgF_{1n1}+hg^{2}F_{2n1}-hF_{1n2}-hgF_{2n2}$
belongs to $S$.
Proof. We have
$A(h,g)_{2}=I_{2}+y_{n}h\begin{bmatrix}g&-1\\ g^{2}&-g\end{bmatrix}$.
$(a)$ Let $\alpha=I-\lambda E_{21},\lambda\in K$. Then
$A(h,g)_{2}^{\alpha}=I_{2}+y_{n}h^{\alpha}\begin{bmatrix}g^{\alpha}&-1\\ (g^{\alpha})^{2}&-g^{\alpha}\end{bmatrix}$
@@
$\alpha_{2}A(h,g)_{2}^{\alpha}\alpha_{2}^{-1}=\begin{bmatrix}1&0\\ -\lambda&1\end{bmatrix}\left(I_{2}+y_{n}h^{\alpha}\begin{bmatrix}g^{\alpha}&-1\\ (g^{\alpha})^{2}&-g^{\alpha}\end{bmatrix}\right)\begin{bmatrix}1&0\\ \lambda&1\end{bmatrix}$
$=I_{2}+y_{n}h^{\alpha}\begin{bmatrix}(g^{\alpha}-\lambda)&-1\\ (g^{\alpha}-\lambda)^{2}&-(g^{\alpha}-\lambda)\end{bmatrix}=A(h^{\alpha},(g^{\alpha}-\lambda))_{2}.$
Consequently, $\alpha A(h,g)^{\alpha}\alpha^{-1}=A(h^{\alpha},(g^{\alpha}-\lambda))$. This means that $A(h,g)\in S$ if and only if $A(h^{\alpha},(g^{\alpha}-\lambda))\in S$. Using this we may assume that $g(0,\ldots,0)=0$.
$(b)$ Set $B=I-fF_{231}\in S$. Then
$B_{3}=\begin{bmatrix}1&0&0\\ -y_{3}f&1&0\\ y_{2}f&0&1\end{bmatrix}$.
@@
$B_{3}A(h,g)_{3}B_{3}^{-1}=I_{3}+y_{n}h\begin{bmatrix}(g-y_{3}f)&-1&0\\ (g-y_{3}f)^{2}&-(g-y_{3}f)&0\\ y_{2}(g-y_{3}f)f&-y_{2}f&0\end{bmatrix}$
[…]
$=A(h,g-y_{3}f)_{3}C_{3},$
where
$C=I+y_{2}h(g-y_{3}f)fF_{3n1}-y_{2}hfF_{3n2}=(I+y_{2}h(g-y_{3}f)fF_{3n1})(I-y_{2}hfF_{3n2}).$
Notice that $C\in S$. Consequently, $A(h,g)\in S$ if and only if $A(h,g-y_{3}f)\in S$. Of course, the same trick can be done with any variable $y_{i}$ if $3\leq i\leq n-1$. Therefore, replacing $g$ by $g-y_{i}f$ for all $3\leq i\leq n-1$ we may assume that
$g=G(y_{1},y_{2})-y_{n}a.$
$(c)$ Now we show how to eliminate $y_{n}a$. First we replace $g$ to
$g+y_{3}a=G(y_{1},y_{2})+y_{3}a-y_{n}a$
using $(b)$. If $\gamma=I+E_{3n}$, then it is easy to check that
$\gamma A(h,g+y_{3}a)^{\gamma}\gamma^{-1}=A(h^{\gamma},(g+y_{3}a)^{\gamma}).$
@@
$(g+y_{3}a)^{\gamma}=g^{\gamma}+(y_{3}+y_{n})a^{\gamma}$
$=G(y_{1},y_{2})-y_{n}a^{\gamma}+(y_{3}+y_{n})a^{\gamma}=G(y_{1},y_{2})+y_{3}a^{\gamma}.$
@@
(d) Let $f$ be a polynomial in $y_{1},y_{2}$. Then $g$ can be replaced by $g+y_{i}f$, where $i=1,2$. For this we first replace $g$ by $g+y_{3}f$ using $(b)$, i.e., we can assume that $g=G(y_{1},y_{2})+y_{3}f$. If $\delta=I+E_{i3}$, then
$\delta A(h,g)^{\delta}\delta^{-1}=(I-h^{\delta}g^{\delta}F_{1n1}-h^{\delta}(g^{\delta})^{2}F_{2n1})A(h^{\delta},g^{\delta}).$
Since $I-h^{\delta}g^{\delta}F_{1n1}-h^{\delta}(g^{\delta})^{2}F_{2n1}\in S$ it follows that $A(h,g)\in S$ if and only if $A(h^{\delta},g^{\delta})\in S$.
Notice that
$g^{\gamma}=G(y_{1},y_{2})+y_{3}f^{\gamma}+y_{i}f^{\gamma}.$
Using $(b)$, we can replace $g^{\gamma}$ by
$G(y_{1},y_{2})+y_{i}f^{\gamma}.$
This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{i}f$ for any $f$. Then we can make $g=0$. Notice that $A(h,0)=I-hF_{1n2}\in S$. $\Box$
Corollary 8.

### v3

@@ 5. Almost tame automorphisms
$\phi=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
$=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
is almost tame. Repeating the proof of Lemma 4 with this data, we get that $D(a)$ is almost tame for all $a$ with $\mathrm{ldeg}(a)\geq 1$. $\Box$
@@
Proof. Consider the decomposition of the exponential automorphism $E([x_{1},x_{2}]a)$ from the proof of Theorem 2. Note that $\mathrm{ldeg}(E([x_{1},x_{2}]a))\geq 4$ implies $\mathrm{ldeg}(a)\geq 1$. Consequently, $D(a)$ is almost tame by Lemma 5. Since all $\phi_{i}$ are Chein automorphisms it follows that $E([x_{1},x_{2}]a)$ is almost tame. Consequently, every exponential automorphism of lower degree $\geq 4$ belongs to $\mathrm{ATAut}(M_{n})$. $\Box$
@@
Every automorphism of the form
$A(h,g)=(x_{1}+[x_{1},x_{n}]hg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]h-[x_{2},x_{n}]hg,x_{3},\ldots,x_{n}),$
where $g,h\in U$, is almost tame.
Proof. $(a)$ Let $\alpha=(x_{1}-\lambda x_{2},x_{2},\ldots,x_{n})$. Then
$A(h,g)^{\alpha}=A(h^{\alpha},g^{\alpha}-\lambda).$
This means that $A(h,g)$ is almost tame if and only if $A(h^{\alpha},g^{\alpha}-\lambda)$ is almost tame. Using this we may assume that $g(0,\ldots,0)=0$.
$(b)$ Consider the almost tame automorphisms
$\phi=(x_{1}-[x_{2},x_{3}]a,x_{2},\ldots,x_{n})$
@@
$\psi=(x_{1}+[x_{3},x_{n}]y_{2}ah(g-y_{3}a),x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{n}]y_{2}ah,x_{3},\ldots,x_{n})$
$=(x_{1}+[x_{3},x_{n}]y_{2}ah(g-y_{3}a),x_{2}-[x_{3},x_{n}]y_{2}ah,x_{3},\ldots,x_{n})$
@@
$A(h,g)^{\phi}=A(h,g-y_{3}a)\psi.$
Consequently, $A(h,g)$ is almost tame if and only if $A(h,g-y_{3}f)$ is almost tame. Of course, the same trick can be done with any variable $y_{i}$ if $3\leq i\leq n-1$. Therefore, replacing $g$ by $g-y_{i}f$ for all $3\leq i\leq n-1$, we may assume that
$g=G(y_{1},y_{2})-y_{n}b,\ \ b\in U.$
$(c)$ Now we show how to eliminate $y_{n}b$. First we replace $g$ to
$g+y_{3}b=G(y_{1},y_{2})+y_{3}b-y_{n}b$
using $(b)$. If $\beta=(x_{1},x_{2},x_{3}+x_{n},x_{4},\ldots,x_{n})$, then it is easy to check that
$A(h,g+y_{3}b)^{\beta}=A(h^{\beta},(g+y_{3}b)^{\beta}).$
@@
$(g+y_{3}b)^{\beta}=g^{\beta}+(y_{3}+y_{n})b^{\beta}$
$=G(y_{1},y_{2})-y_{n}b^{\beta}+(y_{3}+y_{n})b^{\beta}=G(y_{1},y_{2})+y_{3}b^{\beta}.$
@@
(d) Let $c$ be a polynomial in $y_{1},y_{2}$. Then $g$ can be replaced by $g+y_{i}c$, where $i=1,2$. For this, we first replace $g$ by $g+y_{3}c$ using $(b)$, i.e., we can assume that $g=G(y_{1},y_{2})+y_{3}c$. If $\gamma=(x_{1},x_{2},x_{3}+x_{1},x_{4},\ldots,x_{n})$, then
$A(h,g)^{\gamma}=\sigma A(h^{\gamma},g^{\gamma}),$
where
$\sigma=(x_{1},x_{2},x_{3}-[x_{1},x_{n}]h^{\gamma}g^{\gamma}-[x_{2},x_{n}]h^{\gamma}(g^{\gamma})^{2},x_{4},\ldots,x_{n})$
$=(x_{1},x_{2},x_{3}-[x_{1},x_{n}]h^{\gamma}g^{\gamma},x_{4},\ldots,x_{n})(x_{1},x_{2},x_{3}-[x_{2},x_{n}]h^{\gamma}(g^{\gamma})^{2},x_{4},\ldots,x_{n})$
is an almost tame automorphism. Consequently, $A(h,g)$ is almost tame if and only if $A(h^{\gamma},g^{\gamma})$ is almost tame. Note that
$g^{\gamma}=G(y_{1},y_{2})+y_{3}c+y_{1}c.$
Using $(b)$, again we can replace $g^{\gamma}$ by
$G(y_{1},y_{2})+y_{1}c.$
This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{1}c$ for any $c\in K[y_{1},y_{2}]$.
If $\delta=(x_{1},x_{2},x_{3}+x_{2},x_{4},\ldots,x_{n})$ and $g=G(y_{1},y_{2})+y_{3}c$, then
$A(h,g)^{\delta}=\tau A(h^{\delta},g^{\delta}),$
where
$\tau=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]h^{\delta}+[x_{2},x_{n}]h^{\delta}g^{\delta},x_{4},\ldots,x_{n})$
$=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]h^{\delta},x_{4},\ldots,x_{n})(x_{1},x_{2},x_{3}+[x_{2},x_{n}]h^{\delta}g^{\delta},x_{4},\ldots,x_{n})$
is again almost tame automorphism. In this case we have
$g^{\delta}=G(y_{1},y_{2})+y_{3}c+y_{2}c.$
Using $(b)$, we can replace $g^{\delta}$ by
$G(y_{1},y_{2})+y_{2}c.$
This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{2}c$ for any $c\in K[y_{1},y_{2}]$.
Thus we can make $g=0$. Observe that
$A(h,0)=(x_{1},x_{2}-[x_{1},x_{n}]h,x_{3},\ldots,x_{n})$
is a Chein automorphism. $\Box$
Theorem 3.